Midgley Corporation makes a product whose direct labor standards are 2.1 hours per unit and $22.00 per hour. In April, the company produced 6,200 units using 12,780 direct labor-hours. The actual direct labor cost was $269,020.

The labor rate variance for April is:

rev: 11_01_2016_QC_CS-68333

$12,239 U

$12,140 F

$12,239 F

$12,140 U
Answer #1

Solution:

Standard rate of direct labor = $22 per hour

Actual rate of direct labor = $269,020 / 12780 = $21.0500782 per hour

Actual direct labor hours = 12780

Direct labor rate variance = (SR - AR)* AH = ($22 - $21.0500782) * 12780 = $12,140 F